# Noninvasive Acute Compartment Syndrome Diagnosis Using Random Forest Machine Learning

### 18 Jan 2024 v1

@@ Abstract
Acute compartment syndrome (ACS) is an orthopedic emergency, caused by elevated pressure within a muscle compartment, that leads to permanent tissue damage and eventually death. Diagnosis of ACS relies heavily on patient-reported symptoms, a method that is clinically unreliable and often supplemented with invasive intracompartmental pressure measurements. This study proposes a continuous, objective, noninvasive diagnostic for ACS. The device detects ACS through a random forest machine learning model that uses pressure readings from force-sensitive resistors (FSRs) placed on the skin. The final diagnosis is exported real-time to a web application via Bluetooth. To validate the diagnostic, a data set containing FSR measurements and the corresponding simulated intracompartmental pressure was created. The diagnostic achieved an accuracy, on par to the invasive gold standard, of 97%. The device excelled in key performance metrics including precision, sensitivity, and F1 score. Manufactured for 73 USD, our device may be an economic alternative to needle-based diagnostics. These results demonstrate the potential of noninvasive ACS diagnostics to meet clinical standards and enhance patient care.
@@ Introduction
To provide a comprehensive reading of the arm’s skin-to-surface pressure, five Interlink Electronics FSRs were sewn into a polyester biocompatible sleeve. The sensors were wired to an ESP32 Arduino microcontroller, with built-in Bluetooth capability, to enable real-time data transmission and remote monitoring. The microcontroller was placed externally from the sleeve, maximizing patient comfort and preventing device breakage. The device was powered using a USB host device.
@@
As there is a lack of public data sets available on the diagnosis of ACS in relation to skin-to-surface pressure measurements, data points on the raw readings of five FSRs and the true prediction of ACS were gathered. This dataset was collected by placing the diagnostic device on an inflated IV bag to simulate intracompartmental pressure. The IV bag was inflated from zero to 210 mmHg, and the corresponding FSR voltage readings were logged from the Arduino. A total of 200 data points with 20 data points for each pressure of 0, 10, 20, 30, 60, 90, 120, 150, 180 and 210 mmHg were recorded. Simulated pressures greater than 30 mmHg were classified as a true positive prediction.
@@
To classify the detection of ACS using the FSR sensor measurements, a random forest classifier (RFC) machine learning model was deployed to the diagnostic. The model was trained and developed using the python library Scikit-Learn. The RFC algorithm was selected due to its recognized high accuracy in data science and its ability to overcome noise in data (19). The test to train split of the data was 20% to 80%. The input parameters were the raw voltage readings of five FSR sensors and the output was a binary classification of ACS. No feature preprocessing or scaling was performed on the data-set. The RFC was constructed using the standard number of 100 decision trees and all hyperparameters were left at default values. An abbreviated single decision tree from the forest is visualized in Fig. 2, where darker shaded boxes indicate a higher probability for the detection class.
@@
Our model’s performance was assessed using key metrics, including sensitivity, precision, F1 score, accuracy, and specificity. During the training phase, the model achieved a perfect sensitivity. The precision and accuracy during the model’s evaluation were notably high at 0.97. Furthermore, the F1 score reached 0.98, indicating a well-balanced performance between precision and sensitivity. The model also demonstrated an acceptable specificity of 0.88. All these metrics are presented in Fig. 4, where sensitivity, F1 score, precision, accuracy, and specificity were consistently on the higher range from 0 to 1.
Figure 4: Comparison of model performance metrics. Sensitivity, F1 score, precision, accuracy and specificity tended to be of stronger values from the range of 0 to 1.
Out of all metrics, the model exhibited a lower specificity. This suggests a possible challenge in correctly ruling out ACS. The lower specificity may be attributed to a slight imbalance in the training dataset, where fewer instances of the negative class (not ACS) compared to the positive class (ACS) led to a bias in predicting the majority class. Nevertheless, in comparison to prior literature that assessed a diagnostic with similar hardware but a simple average algorithm among pressure readings, our study still proposed a 57% increase in specificity (18).
@@
Moreover, in Fig. 5, a Receiver Operating Characteristic (ROC) curve has been generated to assess the diagnostic performance of the model.
Figure 5: The generated receiver operating characteristic curve of our detection model. The upper-left corner convex shape and related area under curve metric indicates strong performance in distinguishing between true positives and false positives.
The ROC curve visually represents the trade-off between sensitivity and specificity across different threshold settings in the model’s output probability for a binary classification of ACS. The convex shape towards the upper-left corner of our model’s ROC curve suggests a robust ability to correctly classify positive instances while minimizing false positives. This convexity signifies a favorable model performance, as our model outperformed a random classifier (depicted by the dashed gray line) and approached the ideal scenario of perfect classification of the condition.
Altogether, the total manufacturing cost to produce a user ready product was 73 USD. In reference, the industry standard Stryker intracompartmental pressure system is commercially available for 1575 USD (12). Furthermore, our device is 97% cheaper than a 2019 study proposing a noninvasive diagnostic utilizing similar technology that held a total cost of 144.31 USD (18).
In conclusion, our diagnostic model demonstrated exceptional performance metrics, including high precision, accuracy, and F1 score, with a slight trade-off in specificity. The cost-effective manufacturing of our device, priced at 73 USD, makes it a compelling and economic alternative to the current invasive gold standard.
@@ Conclusion
In this study, we developed a noninvasive ACS diagnostic system that utilizes skin-to-surface pressure measurement and RFC machine learning for real-time and continuous detection. To address the lack of public data sets on ACS, our work contributed a dataset containing FSR readings and corresponding simulated intracompartmental pressure. Validation of our diagnostic demonstrated high performance in F1 score, precision, sensitivity, accuracy, and specificity. Notably, our testing suggests that our noninvasive device can achieve accuracy levels comparable to the clinical invasive gold standard and may be able to serve as a cost-effective alternative. Comprehensively, our study presents a promising noninvasive ACS diagnostic that holds the potential to meet clinical standards, become adopted into the standard of care, and improve the patient experience.
@@
As our device was tested using an artificial simulation of intracompartmental pressure, the generalizability of our results to real-world scenarios may be constrained. A continuation of our study could entail in vivo testing to thoroughly validate the efficacy of our diagnostic. Further work including animal models would aid to confirm the device’s accuracy. Additionally, extended tests analyzing the device’s performance in motion-present scenarios have the potential to provide greater insight into the device’s capabilities. Overall, future work is necessary to allow the translation of our device into medical applications.

### 13 Feb 2024 v3

@@ Abstract
Acute compartment syndrome (ACS) is an orthopedic emergency, caused by elevated pressure within a muscle compartment, that leads to permanent tissue damage and eventually death. Diagnosis of ACS relies heavily on patient-reported symptoms, a method that is clinically unreliable and often supplemented with invasive intracompartmental pressure measurements that can malfunction in motion settings. This study proposes an objective and noninvasive diagnostic for ACS. The device detects ACS through a random forest machine learning model that uses surrogate pressure readings from force-sensitive resistors (FSRs) placed on the skin. To validate the diagnostic, a data set containing FSR measurements and the corresponding simulated intracompartmental pressure was created for motion and motionless scenarios. The diagnostic achieved up to 98% accuracy. The device excelled in key performance metrics, including sensitivity and specificity, with a statistically insignificant performance difference in motion present cases. Manufactured for 73 USD, our device may be a cost-effective solution. These results demonstrate the potential of noninvasive ACS diagnostics to meet clinical accuracy standards in real world settings.
@@ Introduction
To provide a comprehensive reading of the arm’s skin-to-surface pressure, five Interlink Electronics FSRs were sewn into a polyester biocompatible sleeve. Each sensor was placed to be in the approximate center of the anterior, lateral, superficial posterior and deep posterior compartments of a simulated lower leg. The sensors were wired to an ESP32 Arduino microcontroller, with built-in Bluetooth capability, to enable real-time data transmission and remote monitoring. The microcontroller was placed externally from the sleeve, maximizing patient comfort and preventing device breakage. The device was powered using a USB host device.
@@
As there is a lack of public data sets available on the diagnosis of ACS in relation to skin-to-surface pressure measurements, data points on the raw readings of five FSRs and the true prediction of ACS were gathered for motion and motionless scenarios. This dataset was collected by placing the diagnostic device on an inflated IV bag to simulate intracompartmental pressure. The IV bag was lined with a 21 mm polyethylene foam layer to simulate subcutaneous fat in the leg and a 2 mm silicone artificial skin. As shown in Fig. 2, the FSRs, sewn in the sleeve, directly touch the artificial skin. The IV bag was inflated from zero to 50 mmHg inclusive, by increments of 10 mmHg, and the corresponding FSR voltage readings were logged from the Arduino. A total of 400 data points with 80 data points for each pressure were recorded. This process was repeated to create another 400 point dataset containing recorded FSR measurements and ground truth pressure in a motion present scenario. Motion was simulated by manually raising and lowering the IV bag, by a 1 m vertical displacement, continuously throughout data collection. This was done to simulate real-life clinical settings such as moving the patient limb, transport conditions, and standing up. Ground truth pressures greater than or equal to 30 mmHg were classified as a positive case. Pressures below 30 mmHg were recorded as a negative ACS case.
Figure 2: Cross-section of testing set-up. Acute compartment syndrome (ACS) was simulated artificially using an intravenous (IV) bag, foam and silicone skin. Ground truth pressure is the inflated pressure of the IV bag.
@@
To classify the detection of ACS using the FSR sensor measurements, a random forest classifier (RFC) machine learning model was deployed to the diagnostic. The model was trained and developed using the python library Scikit-Learn. The RFC algorithm was selected due to its recognized high accuracy in data science, and its ability to overcome noise in data (19). The model was trained using the ideal motionless data set. The test to train split of the data was 20% to 80%. The model was then evaluated using 20% of the motion data set to assess generalization of the model in the alternate use case. The input parameters were the raw voltage readings of five FSR sensors and the output was a binary classification of ACS. No feature preprocessing or scaling was performed on the data-set. The RFC was constructed using the standard number of 100 decision trees and all hyperparameters were left at default values. The max depth of a decision tree was abbreviated to be five to prevent model overfitting of data. This was also done to improve model generalization for a motion present case. An abbreviated single decision tree from the forest is visualized in Fig. 3, where darker shaded boxes indicate a higher probability for the detection class.
@@
Figure 5: Comparison of model performance metrics in motion and no motion present scenarios. Higher scores indicate greater performance. Accuracy, precision, sensitivity, and specificity tended to be of stronger values from the range of 0 to 1 in both test cases. Performance during motionless diagnosis was found to be statistically insignificant compared to the motion present case. Ten percent error bars were used to determine significance.
@@
Our model’s performance in motionless and motion scenarios was assessed using key metrics, including accuracy, precision, sensitivity, specificity and F1 score. The accuracy and precision during the model’s evaluation in the motionless case was notably high at 0.98 and 0.97 respectively. Although accuracy and precision is higher in the motionless case, these metrics are not statistically significantly different from the motion case, when using ten percent error bars. The model in both cases achieved a sensitivity of 0.97, which indicates that fewer cases of ACS are missed. The model also demonstrated a high motionless specificity of 0.98 and an acceptable motion specificity of 0.88. Furthermore, the F1 score reached 0.97 and 0.95 in motionless and motion scenarios, which signals a well-balanced performance between precision and sensitivity. All these metrics are presented in Fig. 5, where accuracy, precision, sensitivity, specificity, and F1 score were consistently on the higher range from 0 to 1 in the two scenarios.
Out of all metrics, the model exhibited a lower specificity in the motion case. This suggests a possible challenge in correctly ruling out ACS. Nevertheless, in comparison to prior literature that assessed a diagnostic with similar hardware but a simple average algorithm among pressure readings, our study still proposed a 57% increase in specificity (18). This is likely because machine learning is more capable of discerning complex relationships between the sensors compared to a simple linear approach.
@@
Moreover, in Fig. 6, a Receiver Operating Characteristic (ROC) curve was generated to assess the diagnostic performance of the model in the motionless scenario. The ROC curve visually represents the trade-off between sensitivity and specificity across different threshold settings in the model’s output probability for a binary classification of ACS. The convex shape towards the upper-left corner of our model’s ROC curve suggests a robust ability to correctly classify positive instances while minimizing false positives.
Figure 6: The generated receiver operating characteristic (ROC) curve of our detection model in the motionless case. The upper-left corner convex shape and related area under curve metric indicates strong performance in distinguishing between true positives and false positives.
This trend is also shown for the ROC curve of the motion scenario shown in Fig. 7. This convexity signifies a favorable model performance, as our model outperformed a random classifier, depicted by the dashed gray line, and approached the ideal scenario of perfect classification of the condition in both test cases. The area under curve (AUC) is greater in the motionless case compared to the motion present scenario. These findings are expected as movement may interfere with FSR measurement and introduce further noise in the data.
Figure 7: The generated receiver operating characteristic (ROC) curve of our detection model in the motion case. The upper-left corner convex shape demonstrates that the classifier’s performance is not due to random chance.
Altogether, the total manufacturing cost to produce a user ready product was 73 USD. In reference, the industry standard Stryker intracompartmental pressure system is commercially available for 1575 USD (12). Furthermore, our device is 97% cheaper than a 2019 study proposing a noninvasive diagnostic utilizing similar technology that held a total cost of 144.31 USD (18). Our device’s cost-effectiveness is most likely due to the use of commercially available FSRs rather than self manufactured ones.
In conclusion, our diagnostic model demonstrated exceptional performance metrics, including high accuracy, precision, sensitivity, and F1 score, with a slight trade-off in specificity in the motion case. The cost-effective manufacturing of our device makes it a compelling and economic alternative to the current invasive gold standard.
@@ Conclusion
In this study, we developed a noninvasive ACS diagnostic system that utilizes skin-to-surface pressure measurement and RFC machine learning for real-time and continuous detection. To address the lack of public data sets on ACS, our work contributed a dataset containing FSR readings and corresponding simulated intracompartmental pressure in both motionless and motion settings. Validation of our diagnostic demonstrated high performance in accuracy, precision, sensitivity, specificity, and F1 score. Notably, our testing suggests that our noninvasive device can achieve accuracy levels comparable to the clinical invasive gold standard and may be able to serve as a cost-effective alternative. Furthermore, our device was able to accurately predict ACS in motion present settings with an insignificant difference compared to motionless detection. This finding demonstrates that a noninvasive diagnostic can be applied to real-world scenarios where continuous pressure monitoring includes periods where the patient is being transported. Additionally, these results indicate that a machine learning noninvasive diagnostic can address past limitations of invasive devices becoming fallible in motion present scenarios (21). Comprehensively, our study presents a promising noninvasive ACS diagnostic that holds the potential to meet clinical standards, become adopted into the standard of care and improve the patient experience.
@@
As our device was tested using an artificial simulation of intracompartmental pressure, the generalizability of our results to real-world scenarios may be constrained. A continuation of our study could entail in vivo testing to thoroughly validate the efficacy of our diagnostic. Further work including animal models would aid to confirm the device’s accuracy. Additionally, extended tests analyzing the device’s performance in long-term continuous applications have the potential to provide greater insight into the device’s capabilities. Overall, future work is necessary to allow the translation of our device into medical applications.